# DELTA: Dense Efficient Long-range 3D Tracking for any video

### v1

@@ Appendix A Appendix
Qualitative results of long-range 2D flows are visualized in 6. Thanks to the strong temporal consistency of our method, it produces smooth predictions over time, effectively avoiding the ’flicker’ artifacts that are commonly observed in per-frame optical flow predictions, such as those produced by DOT(Le Moing et al., 2024).
We also compare our long-range flow predictions with the current SoTA optical flow model, SEA-RAFT (Wang et al., 2025), and DOT on in-the-wild videos. The results are visualized in figure 7 showing that our approach achieves better temporal smoothness and less artifact for long-range motion prediction.
@@
Dynamic video pose estimation. Our proposed approach efficiently achieves dense 3D tracking across multiple keyframes throughout the video. As an illustration, consider the dense 3D trajectories $\mathcal{P}^{(0,)}\in\mathbb{R}^{T\times H\times W\times 4}$ of pixels in the first frame. For any destination frame $\nu$ the 3D points $\mathcal{P}^{(0,\nu)}=\mathcal{P}^{(0,)}[\nu]\in\mathbb{R}^{H\times W\times 4}$ represent the 3D positions of all pixels of frame 0 in the camera coordinate at $\nu$, naturally combining camera motion from 0 to $\nu$ along with any non-rigid scene motion within that interval. This setup aligns with MonST3R (Zhang et al., 2024), a dynamic version of DUSt3R (Wang et al., 2024b). Consequently, we apply the global alignment approach used in Wang et al. (2024b); Zhang et al. (2024) to estimate camera poses across dynamic video sequences.
@@
In detail, we uniformly sample keyframes with a stride of 2 from the input video and densely track all pixels of these frames. For each keyframe $k$, tracking is performed both forward and backward within a window $[k-w,k+w]$, where $w=8$. Following Zhang et al. (2024), we model the pose estimation as an optimization task, where the learnable parameters include per-keyframe depth maps, as well as per-keyframe intrinsic and extrinsic camera parameters, and optimize with gradient descent. The objective function, similar to Zhang et al. (2024), combines alignment loss, temporal smoothness loss, and 2D flow loss. For the 2D flow loss, we calculate pseudo optical flow ground truth using our dense tracking approach rather than relying on an off-the-shelf model (Wang et al., 2025; Teed & Deng, 2020). For non-keyframe frames, camera poses are interpolated from the two nearest neighboring keyframes.
@@
We evaluate the pose estimation performance on multiple datasets, including Sintel (Butler et al., 2012) and TUM-dynamics (Sturm et al., 2012). On Sintel, we follow the evaluation protocol in Zhang et al. (2024); Chen et al. (2024), which excludes static and easy scene, remaining 14 test sequences. For TUM-dynamics, we sample the first 90 frames with the temporal stride of 3. We report the Absolute Translation Error (ATE), Relative Translation Error (RPE trans), and Relative Rotation Error (RPE rot), after applying a Sim(3) Umeyama alignment on prediction to the ground truth. The results are reported in Table 8.
@@
Ours 0.195 0.075 0.549 0.052 0.007 1.343
Table 8: Pose estimation results on Sintel and TUM datasets. Our method achieves competitive results compared to other approaches specifically designed for visual odometry or SLAM tasks.

### v3

@@ A.1 Implementation details
A.4 More results
Qualitative results of long-range 2D flows are visualized in figure 6. Thanks to the strong temporal consistency of our method, it produces smooth predictions over time, effectively avoiding the ’flicker’ artifacts that are commonly observed in per-frame optical flow predictions, such as those produced by DOT(Le Moing et al., 2024).
Figure 7: Comparison of long-range optical flow predictions on Sintel dataset between our approach (with CNN-based/attention-based upsampler) and SEA-RAFT. We predict optical flows from the first frame to subsequent frames of the video. While SEA-RAFT excels at short-range flow prediction (first column), it fails to predict flow between far-away frames. In contrast, our approach, with an attention-based upsampler, achieves smooth and consistent predictions across frames, outperforming the baseline using convolutional upsampler.
CVO (Extended)
EPE $\downarrow$ OA $\uparrow$
W/o anchor tracks 4.50 / 3.10 / 6.95 69.7
With anchor tracks 3.63 / 2.67 / 5.24 71.6
Table 9: Ablation on the role of anchor tracks.
Qualitative comparison of the design of upsampler on the Sintel dataset (Butler et al., 2012) is visualized in figure 7. Our attention-based upsampler (last row) significantly outperforms the CNN-based upsampler from RAFT (Teed & Deng, 2020) (second row), with better detail and less artifacts. For reference, we include predictions from the state-of-the-art optical flow model SEA-RAFT (Wang et al., 2025) (first row), which, despite being trained on Sintel, excels at short-range predictions (first column) but fails with long-range flows.
We also compare our long-range flow predictions between SEA-RAFT, DOT, and our method on in-the-wild videos. The results are visualized in figure 8 showing that our approach achieves better temporal smoothness and less artifact for long-range motion prediction.
@@ A.4 More results
Ablation on the anchor tracks is shown in Table 9. Removing the anchor tracks during training increases the EPE by nearly 1.0, meaning that the anchor tracks are important to avoid train-test mismatch.
A.5 Further discussions on the 3D tracking performance
As shown in Table 5 of the main paper, baseline methods (2D tracking + depth estimation) significantly underperform compared to our end-to-end 3D tracking approach. Interestingly, better 2D tracking models do not always lead to better 3D tracking performance (e.g., BootsTAPIR versus CoTracker). This is because the performance of these baselines heavily relies on the quality and scale-consistency of the input video depth. However, frame-wise depth estimators like ZoeDepth and UniDepth lack scale-consistency (see our webpage), where even small depth errors can cause significant 3D location discrepancies.
Furthermore, 2D tracking + depth estimation can only predict motion for visible points within the viewpoint, restricting applications such as 3D/4D reconstruction, which require complete motion predictions, including occluded and out-of-frame regions.
These challenges underline the need for end-to-end 3D tracking models to achieve reliable 3D performance. Our approach addresses these limitations and is also complementary to recent advances in 2D point tracking, enabling us to integrate their findings to further enhance both 2D and 3D tracking performance.
A.6 Application
Dynamic video pose estimation. DUSt3R (Wang et al., 2024b) introduced a paradigm for estimating static 3D scene geometry and camera poses from image sets. By training a model on large-scale data, it predicts aligned 3D point maps for image pairs, followed by lightweight optimization to obtain scale-consistent depth maps and camera poses. However, this approach fails on videos with dynamic objects. MonST3R (Zhang et al., 2024) extended DUSt3R to dynamic videos by fine-tuning the model on dynamic datasets. Our method aligns with this approach.
@@ A.6 Application
We first perform dense pixel tracking from the initial frame, producing dense 3D trajectories $\mathcal{P}^{(0,)}\in\mathbb{R}^{T\times H\times W\times 4}$. For any destination frame $\nu$, the 3D points $\mathcal{P}^{(0,\nu)}=\mathcal{P}^{(0,)}[\nu]\in\mathbb{R}^{H\times W\times 4}$ represent the 3D positions of all pixels from the initial frame in the camera coordinate system at frame $\nu$. This naturally combines the camera motion from frame 0 to $\nu$ with any non-rigid scene motion over the interval. This setup aligns with the methodology of MonST3R (Zhang et al., 2024). We then apply the global alignment approach from Wang et al. (2024b); Zhang et al. (2024) to estimate camera poses across dynamic video sequences.
@@
In detail, we uniformly sample keyframes with a stride of 2 from the input video and densely track all pixels of these frames. For each keyframe $k$, tracking is performed both forward and backward within a window $[k-w,k+w]$, where $w=8$. Following Zhang et al. (2024), we model the pose estimation as an optimization task, where the learnable parameters include per-keyframe depth maps, as well as per-keyframe intrinsic and extrinsic camera parameters, and optimize with gradient descent. The objective function combines alignment loss, temporal smoothness loss, and 2D flow loss. For the 2D flow loss, we calculate pseudo optical flow ground truth using our dense tracking approach rather than relying on an off-the-shelf model (Wang et al., 2025; Teed & Deng, 2020). For non-keyframe frames, camera poses are interpolated from the two nearest neighboring keyframes.
@@
We evaluate the pose estimation performance on Sintel (Butler et al., 2012) and TUM-dynamics (Sturm et al., 2012) datasets. On Sintel, we follow the evaluation protocol in Zhang et al. (2024); Chen et al. (2024), which excludes static and easy scene, remaining 14 test sequences. For TUM-dynamics, we sample the first 90 frames with the temporal stride of 3. We report the Absolute Translation Error (ATE), Relative Translation Error (RPE trans), and Relative Rotation Error (RPE rot), after applying a Sim(3) Umeyama alignment on prediction to the ground truth. The results are reported in Table 10. Our method demonstrates comparable performance to state-of-the-art approaches specifically tailored for visual odometry and SLAM tasks. We visualize the camera pose and dynamic 3D scene reconstruction on casual videos in figure 9.
@@
Ours 0.172 0.060 0.553 0.052 0.007 1.343
Table 10: Pose estimation results on Sintel and TUM datasets. The upper group (first four rows) includes methods that estimate camera poses only, without reconstructing scene geometry while the lower group and our approach provide both camera poses and per-frame depth maps. Our method achieves competitive results compared to other approaches specifically designed for visual odometry or SLAM tasks.